# IEEEexample:BSTcontrol

### v3

@@ 3.2.2 Modeling the SPECT system
We simulated the scenario where the torso region of a patient was imaged for 30 minutes using a dual-head GE Discovery 670 SPECT/CT system equipped with NaI(TI) crystals and a HEGP collimator [45]. The simulations were performed with SIMIND. The intrinsic spatial and energy resolution of the scintillation detector in the system were 3.9 mm and 9.8 % at 140 keV, respectively. Projections were acquired in four energy windows, as indicated in Fig. 1, and at 60 angular positions spaced uniformly over ${}^{\circ}$. We simulated the emission spectra of both isotopes, the stray-radiation-related noise, and all relevant image-degrading processes in SPECT. In the following sections, each instance of SPECT projection generated for a simulated patient is referred to as a noise realization.
@@ 3.3.1 Evaluating the convergence of the MEW-PDQ method
[…]
@@ 5 Discussion
We have designed, implemented, and evaluated a MEW-PDQ method to jointly quantify the activity uptake of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ in $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs. The method was designed to address the challenges in quantitative SPECT for $\alpha$-RPTs, including the extremely low counts, crosstalk contaminations between $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$, and the complicated isotope and SPECT physics. The results (Figs. 5-12) demonstrate the efficacy and superiority of the method over conventional RBQ methods in jointly estimating the regional uptake of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ in multiple scenarios.
@@
The results in Figs. 5-7 demonstrate that the MEW-PDQ method yielded accurate estimates of lesion uptake of both $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ with varying lesion characteristics, including different lesion sizes (Fig. 5 a), LBURs (Fig. 6 a), and $\mathrm{{}^{227}Th}$ to $\mathrm{{}^{223}Ra}$ uptake ratios (Fig. 7 a) in the case where there was homogeneous uptake in the VOIs. We note that this method, instead of estimating a large number of voxel values (as is done by the conventional RBQ methods) from projections with very low counts, is estimating uptake in a smaller number of VOIs. Posing the problem this way helps the method achieve close to zero bias. In contrast, as is well known, ML-based pixel value quantification [53] or conventional RBQ methods that employ ML-based image reconstruction [54] can yield biased estimates at low counts, as observed in this study as well. The low bias with the proposed method also illustrates the efficacy of the method in compensating for image-degrading effects, including crosstalk contamination, even at low count levels. These observations highlight the significance of the proposed method for the task of estimating mean uptake for $\mathrm{{}^{227}Th}$-based RPTs. Further, the proposed method was observed to be relatively insensitive to PVEs, as demonstrated by the high accuracy for even small lesion sizes. We note that the DOSEM method is specifically designed to account for the crosstalk contamination, and the GTM-based method is specifically designed to compensate for PVEs. However, the proposed method outperformed both approaches.
In addition, the proposed method is also highly precise and yielded NSD values close to those derived from CRLBs (Figs. 5-7 b and 9), indicating that the method had the theoretically highest possible precision as an unbiased estimator. This high precision of the method is especially significant given the low counts in $\alpha$-RPT SPECT. Note that the proposed method is an ML estimator. If an efficient estimator exists, the ML estimator is efficient [30]. Thus, these results indicate that the MEW-PDQ method may be an efficient estimator.
@@
An important feature of the MEW-PDQ method is the use of multiple energy windows in quantification. As theoretically shown through the application of the data-processing inequality, discarding projections in one or more energy windows decreases the Fisher information for estimating parameters from the acquired data [35]. This theoretical finding is manifested in our results in Fig. 8, where we see that using data from multiple energy windows reduces the CRLB significantly. This further motivates the need to develop quantitative SPECT methods that can use data from multiple energy windows, such as the proposed method. This is particularly crucial for $\alpha$-particle SPECT, as the low number of detected counts can be increased by including data from multiple energy windows. Further, as shown in Fig. 9, using projections in all four energy windows, the variance of estimates obtained by the proposed method approached the corresponding CRLB for all VOIs and both isotopes. This finding demonstrates the effectiveness of the proposed method in extracting information from data contained in these energy windows. Overall, these results show that effectively using data from multiple energy windows in jointly quantifying regional uptake of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$, as can be enabled by the proposed method, can yield improved quantification.
@@
The proposed MEW-PDQ method does not explicitly account for spatial intra-regional uptake heterogeneity, leading to model mismatches in scenarios where such heterogeneity is present. Fig. 11 shows that as intra-regional uptake heterogeneities in larger regions increase, both the absolute NB and NRMSE of the estimated lesion uptake from the proposed method increase as well. However, with moderate levels of intra-regional heterogeneities, such as for the patients simulated by setting $\mathcal{W}$ to 0.4 or lower in Eq. (24), the proposed method still consistently outperforms the conventional RBQ methods. This finding suggests that the MEW-PDQ method could still be a reliable choice with better performance than conventional RBQ methods for moderate levels of intra-regional heterogeneity. It is essential to note, however, that the actual extent of intra-regional heterogeneity in patients undergoing $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs remains unclear. The range of intra-regional uptake heterogeneities evaluated in this study was deliberately extensive to ensure a comprehensive understanding of the methods in various conditions, including those that pose significant challenges. Additionally, to leverage the reliable performance of the proposed method in clinical applications, a practical strategy is to carefully delineate VOIs using both the CT and reconstructed SPECT images to consider regions that have low intra-regional heterogeneities.
Fig. 12 illustrates that inaccuracies in VOI definitions can affect the accuracy of the estimates produced by both the proposed and conventional RBQ methods. As the degree of inaccuracy in VOI definition due to rigid-body transformation increases, the accuracy of all the considered methods decrease. However, at each degree of inaccuracy, the proposed MEW-PDQ method consistently outperformed the conventional RBQ methods. Additionally, the results in Fig. 12 show that, compared with the conventional RBQ methods, the proposed method also exhibited lower sensitivity to inaccuracy in VOIs arising from non-rigid transformations caused by patient breathing during the SPECT scanning process.
@@
A tradeoff with the proposed method is the ability to estimate regional uptake rather than individual voxel values. The estimated regional uptake provide inputs for organ and lesion dosimetry, which is often is the clinical task of interest [10, 12]. However, this then precludes the estimation of dose-volume histograms for VOIs such as lesions. However, the low count rate typical for $\alpha$-RPT SPECT leads to reconstructed images that are erroneous for quantification tasks. For example, as observed in our study, when using the DOSEM-based method, the NSD of estimated voxel values for $\mathrm{{}^{223}Ra}$ within the center of the lesion region were as high as 300%. This can thus lead to highly erroneous dose-volume histograms even with a conventional RBQ method. Consequently, a more viable objective is to estimate regional uptake. For this task of estimating regional uptake, we observe across multiple scenarios in this manuscript that the MEW-PDQ method is reliable and significantly outperforms the conventional RBQ methods.
@@ Acknowledgment
[…]

### v4

@@ 3.2.2 Modeling the SPECT system
We simulated the scenario where the torso region of a patient was imaged for 30 minutes using a dual-head GE Discovery 670 SPECT/CT system equipped with NaI(TI) crystals and a HEGP collimator [45]. The simulations were performed with SIMIND. The intrinsic spatial and energy resolution of the scintillation detector in the system were 3.9 mm and 9.8 % at 140 keV, respectively. Projections were acquired in four energy windows, as indicated in Fig. 1, and at 60 angular positions spaced uniformly over 360. We simulated the emission spectra of both isotopes, the stray-radiation-related noise, and all relevant image-degrading processes in SPECT. In the following sections, each instance of SPECT projection generated for a simulated patient is referred to as a noise realization.
@@ 3.3.1 Evaluating the convergence of the MEW-PDQ method
In this experiment, the simulated patients had the same height and weight as a 50th percentile male adult in the US [44], which we refer to as average patient size. The patient torsos contained a total of 11 MBq activity of both $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$, referred to as standard total uptake. The regional uptake ratios of the two isotopes in the four VOIs of the patients are shown in Table 1. Based on the analysis of the clinical SPECT of patients receiving $\mathrm{{}^{223}Ra}$ therapy, these values were chosen to simulate the differences in uptake in different VOIs. The convergence of the proposed method was evaluated for lesions with different diameter values. We simulated three patients, each had a lesion with diameters: 15 mm, 25 mm, and 35 mm, respectively. We generated one noise realization for each patient, as described in Sec. 3.2.2, then applied the MEW-PDQ method for 1500 iterations, and computed the error in the estimated lesion uptake for both isotopes after every 50 iterations for each patient.
@@ 5 Discussion
We have designed, implemented, and evaluated a MEW-PDQ method to jointly quantify the activity uptake of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ in $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs. The method was designed to address the challenges in quantitative SPECT for $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs, including the extremely low counts, crosstalk contaminations between $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$, and the complicated isotope and SPECT physics. The results (Figs. 5-12) demonstrate the efficacy of the method in jointly estimating the regional uptake of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ in multiple scenarios.
@@
The results in Figs. 5-7 demonstrate that the MEW-PDQ method yielded accurate estimates of lesion uptake of both $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ and outperformed the DOSEM and GTM-based methods with varying lesion characteristics, including different lesion sizes (Fig. 5 a), LBURs (Fig. 6 a), and $\mathrm{{}^{227}Th}$ to $\mathrm{{}^{223}Ra}$ uptake ratios (Fig. 7 a) in the case where there was homogeneous uptake in the VOIs, consistent with the assumption of the proposed method. The close-to-zero bias estimates for both isotopes highlight the efficacy of the proposed method in accounting for the crosstalk contamination, one of the major image-degrading effects in quantitative SPECT for $\mathrm{{}^{227}Th}$-based RPTs. Further, the proposed method was observed to be relatively insensitive to PVEs, as demonstrated by the high accuracy for even small lesion sizes. We note that the proposed method performs direct ML estimation of the targeted VOI uptake, and ML estimation is known to be asymptotically unbiased. In addition, the proposed method yielded NSD values close to those derived from CRLBs (Figs. 5-7 b and 9). Again, the proposed method is an ML estimator. If an efficient estimator exists, the ML estimator is efficient [30]. Thus, these results indicate that the MEW-PDQ method may be an efficient estimator.
@@
An important feature of the MEW-PDQ method is the use of multiple energy windows in quantification. As theoretically shown through the application of the data-processing inequality, discarding projections in one or more energy windows decreases the Fisher information for estimating parameters from the acquired data [35]. This theoretical finding is manifested in our results in Fig. 8, where we see that using data from multiple energy windows reduces the CRLB significantly. This further motivates the need to develop quantitative SPECT methods that can use data from multiple energy windows, such as the proposed method. This is particularly important for $\alpha$-particle SPECT, as the low number of detected counts can be increased by including data from multiple energy windows. Further, as shown in Fig. 9, using projections in all four energy windows, the variance of estimates obtained by the proposed method approached the corresponding CRLB for all VOIs and both isotopes. This finding demonstrates the effectiveness of the proposed method in extracting information from data contained in these energy windows. Overall, these results show that effectively using data from multiple energy windows in jointly quantifying regional uptake of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$, as can be enabled by the proposed method, can yield improved quantification.
@@
The proposed MEW-PDQ method does not account for spatial intra-regional uptake heterogeneity, leading to model mismatches in scenarios where such heterogeneity is present. As illustrated in Fig. 11, as the level of intra-regional uptake heterogeneity increases, the absolute NB and NRMSE of the estimated lesion uptake from the proposed method also increase. Despite this, with moderate levels of intra-regional heterogeneities, the proposed method still consistently outperforms the conventional RBQ methods. This finding suggests that the MEW-PDQ method could still be a reliable choice with better performance than conventional RBQ methods given such moderate levels of intra-regional heterogeneity. However, given high levels of intra-regional heterogeneity, the proposed method may yield less reliable estimates of regional uptake compared to the conventional RBQ methods (Fig. 11). Under those conditions, the model mismatch for the proposed method in representing the activity uptake distribution within VOIs may impact the performance of the method. We note that the actual extent of intra-regional heterogeneity in patients undergoing $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs remains unclear. The range of intra-regional uptake heterogeneities evaluated in this study was deliberately extensive to ensure a comprehensive understanding of the methods in various conditions, including those that pose significant challenges. Additionally, to leverage the reliable performance of the proposed method in clinical applications, a practical strategy is to carefully delineate VOIs using both CT and reconstructed SPECT images to define VOIs with low intra-regional heterogeneities. However, not accounting for intra-regional heterogeneity remains a limitation of the proposed method when defining VOIs with relatively homogeneous uptake is not achievable. To address this limitation, a Wiener estimator-based approach is being developed to estimate regional uptake from SPECT projections while accounting for intra-regional heterogeneity [53]. Extending the proposed method to model intra-regional heterogeneity is an important area of future research.
Fig. 12 illustrates that inaccuracies in VOI definitions can affect the accuracy of the estimates produced by both the proposed and conventional RBQ methods. In these results, as the degree of inaccuracy in VOI definition due to rigid-body transformation increased, the accuracy of all the considered methods decreased. However, at each degree of inaccuracy, the proposed MEW-PDQ method consistently outperformed the conventional RBQ methods. Additionally, these results show that, compared with the conventional RBQ methods, the proposed method also exhibited lower sensitivity to inaccuracy in VOIs arising from non-rigid transformations caused by patient breathing during the SPECT scanning process.
@@
The proposed method is designed to directly estimate uptake within volumes of interest, thus providing inputs for organ and lesion dosimetry, which is often the clinical task of interest [10, 12]. However, since the method does not yield a voxel-based reconstruction, it cannot be used for applications that require such reconstructions, such as voxel-based dosimetry. For these applications, there remains an important need to develop voxel-based reconstruction methods for $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs, and this is an important area of future research. For example, if it is known that the VOIs do not have fully homogeneous uptake but have relatively homogeneous uptake in some VOIs and highly heterogeneous uptake in other VOIs, reconstruction methods could represent the VOIs using a small number of voxels with non-uniform sizes, similar to those proposed in previous studies [54, 55]. Similar to the proposed method, such methods could make the estimation problem less ill-posed and reduce the noise in the estimated voxel values. Additionally, more recently, voxel-based reconstruction methods that use list-mode data and avoid binning-related information loss have been developed for single-isotope $\alpha$-RPTs [34]. Reconstructions from list mode data could extract the maximum possible information from the limited number of detected photons, thereby enhancing the quality of reconstructed images. Overall, developing and validating advanced voxel-based image reconstruction methods for $\mathrm{{}^{227}Th}$-based $\alpha$-RPTs is an important area of research; future studies could also compare the performance of the proposed method with RBQ methods that are based on such reconstructions.
@@ Acknowledgment
To compare simulated and physically acquired projection data, we used a NEMA phantom (Data SpectrumTM, USA). The spheres of this phantom were filled with $\mathrm{{}^{227}Th}$ solutions with an activity concentration of 40 kBq/ml. The rest of the phantom was filled with water to simulate attenuation and scatter due to soft tissue. The phantom was scanned on a GE Discovery 670 SPECT/CT system with a medium energy general purpose (MEGP) collimator one day after purifying the $\mathrm{{}^{227}Th}$ isotope and filling the isotope in the phantom. Thus, at the time of scanning, a small portion of the $\mathrm{{}^{227}Th}$ had decayed to $\mathrm{{}^{223}Ra}$. During imaging, projections were acquired in two energy windows, corresponding to the two major photopeaks of $\mathrm{{}^{223}Ra}$ (66 - 98 keV) and $\mathrm{{}^{227}Th}$ (217 - 260 keV) at 60 angular positions spaced uniformly over 360. The same image-acquisition process was modeled using our simulation approach (Sec. III B in the main manuscript). In the simulation, the concentration of $\mathrm{{}^{227}Th}$ and $\mathrm{{}^{223}Ra}$ were calculated theoretically, based on the filling and scanning time. The profiles of the projection data obtained with the physical scanner and with the simulation approach from the two energy windows were compared.